Saif is planning to buy a house in 19 years. He wants to invest RO 334 now and hopes to have RO 9334 to spend on the house when he buys it. What kind of interest rate would he need if his investment is compounded monthly?
Answer by Boreal(15235)

9334=334+(1+(r/12))^228
27.946=(1+(r/12))^228
don't round
ln both sides
3.330=228 ln(1+(r/12))
0.01460=ln(1+(r/12))
raise to e power
1.0147=1+(r/12)
0.0147=r/12
r=17.656%
--
this is doubling more than 4 times, and rule of 72 suggests 4 year doubling time, so this doubles four times and almost a 5th
334*2^4=5344, and one more doubling would be over 9334.
alternatively, ln 27=3.33, so in 19 years needs an interest rate of 333/19 in percent, and that is 17.53%, so answer is reasonable.